The expression f(x+h)-f(x) over h, is called the difference quotient. Determine the difference quotient of f(x)=2x^2-4x.
Anybody know the formula for this or what the first step is at least? Thanks!

Answer by Alan3354(69443)

The expression f(x+h)-f(x) over h, is called the difference quotient. Determine the difference quotient of f(x)=2x^2-4x.
----------------
Find f(x+h) by subbing x+h for x:
f(x+h) = 2(x+h)^2 - 4(x+h)
= 2(x^2 + 2xh + h^2) - 4x - 4h
f(x+h)= 2x^2 + 4xh + 2h^2 -4x - 4h
f(x) = 2x^2 - 4x
------------------ Subtract
dq*h = 2h^2 + 4xh - 4h
Divide by h:
= 4x - 4 + 2h
----------------